OpenStudy (anonymous):

any who helps I will give a medal! 1. Provide a counterexample that shows the statement is false. Explain why the counterexample makes the statement false. If two fractions have unlike denominators, then the LCD is the product of their denominators

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you know what you are being asked to do?

OpenStudy (anonymous):

LCD of a/b and c/d is given by LCD(a,c)/LCM(b,d). Consider example, 5/6 and 4/15. Here LCD=LCD(5,4)/LCM(6,15)=1/30. However, the product of their denominators is 60. This is a counter example.

OpenStudy (anonymous):

LCD of a/b and c/d is given by LCD(a,c)/LCM(b,d). Consider example, 5/6 and 4/15. Here LCD=LCD(5,4)/LCM(6,15)=1/30. However, the product of their denominators is 60. This is a counter example.

OpenStudy (anonymous):

no can you help me @misty1212

OpenStudy (misty1212):

ok sure

OpenStudy (misty1212):

If two fractions have unlike denominators, then the LCD is the product of their denominators you want a "counter example" that means you want an example of two fractions with unlike denominators where the least common multiple of the denominators is NOT their produce

OpenStudy (misty1212):

so for example if the denominators had a common factor, like 4 and 12, then the least common multiple of 4 and 12 is just 12, not \(4\times 12\) that is all you need for a "counter example"

OpenStudy (misty1212):

in simple english you have to provide an example of two fractions whose least common denominator is NOT found by multiplying the denominators together

OpenStudy (anonymous):

ok